# SceneMI: Motion In-betweening for Modeling Human-Scene Interactions

### v1

@@ Abstract
**footnotetext: Work done during an internship at Snap Research NYC, Snap Inc.$\dagger$\$$\dagger$\$footnotetext: Co-corresponding author
@@ 1 Introduction
In this work, we propose SceneMI, a conditional diffusion model tailored for scene-aware motion in-betweening. We extract hierarchical scene features: the global scene is represented by a coarse occupancy voxel grid, while local scenes are encoded as Basis Point Set (BPS) features computed at keyframe time steps based on selected surface body markers. This scene encoding effectively captures the high-level scene layout and spatial context anchored on immediate keyframes, preventing overfitting on specific local scene geometries. We apply imputation to keyframes to effectively encode the temporal location of keyframes within the motion sequence. We particularly leverage the inherent denoising capabilities of diffusion models to handle noisy keyframes. During inference, noisy keyframes guide the denoising sampling from steps $T$ until $T^{*}$ (a hyperparameter). For the remaining steps (from $T^{*}-1$ to $1$), the full motion sequences, including noisy key poses, are iteratively denoised.
@@ 2.1 Human Motion Synthesis
Advancements in data-driven approaches have led to appealing breakthroughs in human motion synthesis. Beyond achieving high-quality and realistic motion outputs, many applications are designed to generate motions adaptable to various conditions, such as action categories [15, 55, 68, 8, 24], text [4, 17, 18, 85, 84, 16, 68, 88, 56, 43, 57], audio [35, 14, 1, 65, 69, 41, 60], interacting objects [80, 12, 40, 92, 6, 34, 67, 3, 39], and scenes [66, 28, 75, 27, 42, 36, 93, 83, 7, 25, 76, 48, 47]. Recently, there has been growing interest in trajectory-based controllable motion generation. For instance, PriorMDM [64] fine-tuned an existing motion diffusion model MDM [68] to support end-effector trajectory control. GMD [64] incorporated classifier-free guidance into a diffusion model to enable root joint-based controllable generation. OmniControl [78] expanded this capability to arbitrary joint trajectories using a combination of ControlNet [87] and inference-time guidance. The recent ControlMM [58] approach, based on masked motion modeling, further enables multi-joint trajectory-based controllable motion generation.
@@ 3.1 Scene Encoding
The global scene features represent the overall spatial context of the entire scene as an occupancy voxel grid at a coarse resolution of $\mathbf{c}_{g}\in\{0,1\}^{d_{x}\times d_{y}\times d_{z}}$ with 0.1m per voxel, where a value of 1 indicates an occupied voxel and 0 represents an unoccupied space. The center and orientation of this grid are initialized by the root position and orientation in the first frame. To effectively encode these global scene features, we employ a Vision Transformer (ViT) model, following the approach in [28]. This global scene encoder processes inputs with dimensions $48\times 24\times 48$ and results in a 512-dimensional feature vector. The global scene encoder is jointly trained with the rest of the framework.
@@ 3.2 Scene-aware Motion In-betweening
The motion-related features are obtained by concatenating the features of $\tilde{\mathbf{x}}_{t}=\texttt{\small spatialconcat}(\mathbf{x}_{t}^{\prime},\mathbf{c}_{l}^{\prime},\mathbf{m})$. These features are further concatenated with body shape features $\mathbf{b}$ and global scene features $\mathbf{c}_{g}$ along the temporal dimension. Subsequently, the embeddings of diffusion steps (i.e., $t$) are added to all input features, as illustrated in Fig. 2. During training, we randomly mask out $\mathbf{c}_{g}$ with a probability of $10\%$ to enable classifier-free guidance during inference.
@@ 3.2.1 Handling Noisy Keyframes
In real-world scenarios, when the keyframe poses are captured by less accurate mocap sensors or extracted from videos, directly using noisy keyposes for motion inbetweening would compromise the output motion quality. Fortunately, diffusion models inherently add noise to motion data during the forward process, and the models learn to remove this noise during the reverse process. As suggested by MotionMix [24] and Rohm [89], the noise in imperfect motions could be analogous to the added noise in the diffusion process at certain timesteps. Motivated by this insight, we divide the diffusion/sampling timesteps into two ranges: $[T,T^{*}+1]$ and $[T^{*},1]$. From early diffusion timestep $T$ to $T^{*}+1$, the prediction of $\mathbf{x}_{0}$ is guided by noisy keyposes $\mathbf{s}^{\text{noisy}}$ for keyframe alignment, while for the remaining timesteps ($T^{*}$ to $1$), both keyframes and in-betweening frames are iteratively denoised by $\mathcal{D}_{\theta}$.
@@ 4 Experiments
For training, we utilized TRUMANS [28], the largest high-quality motion capture dataset (30 FPS) with hand-crafted scene geometry. Motion sequences were segmented into $N=121$ frames, with 20% randomly selected for the test set. To train for noisy keyframe handling, we added zero-mean noise with a standard deviation determined by noise level $l$, defined as $(l^{\circ},l^{\circ},l\ \text{cm})$for the ground-truth SMPL parameters$(\phi,\psi,\gamma)$. For evaluation, we tested our approach on both noise-free and synthetic noisy TRUMANS test sets. To further assess out-of-domain generalization robustness, we evaluated performance on the real-world GIMO dataset [95] and demonstrated video-based HSI reconstruction using the PROX dataset [21].
@@ 4.1 Evaluation on Noise-free TRUMANS
We compared our model to ablated variants that exclude global scene features $\mathbf{c}_{g}$, local scene features $\mathbf{c}_{l}$, or both (i.e.,“w/o scene awareness“). We also evaluated performance on close-proximity interaction frames where the minimal distance between human and scene is less than 8 cm, with results shown in Tab. 2. From Tabs. 2 and 1, we observed that scene-awareness effectively reduces the collision ratio from 0.237 to 0.162 in interaction frames and enhances alignment with ground-truth motions by over 50% (MJPE from 0.059 to 0.023 m). This observation is consistent with the visualized generation results in Fig. 3. Both global and local scene features enhance motion quality and alignment, with global features playing more prominent roles. Refer to the supplementary video for more results.

### v2

@@ Abstract
**footnotetext: Most work done during an internship at Snap Research NYC, Snap Inc.$\dagger$\$$\dagger$\$footnotetext: Co-corresponding author
@@ 1 Introduction
In this work, we propose SceneMI, a conditional diffusion model tailored for scene-aware motion in-betweening. We extract hierarchical scene features: the global scene is represented by a coarse occupancy voxel grid, while local scenes are encoded as Basis Point Set (BPS) features computed at keyframe time steps based on selected surface body markers. This scene encoding effectively captures the high-level scene layout and spatial context anchored on immediate keyframes, preventing overfitting on specific local scene geometries. We apply imputation to keyframes to effectively encode the temporal location of keyframes within the motion sequence. We particularly leverage the inherent denoising capabilities of diffusion models to handle noisy keyframes. During inference, noisy keyframes guide the denoising sampling from steps $T$ until $T^{*}$ (a hyperparameter). For the remaining steps (from $T^{*}-1$ to 111), the full motion sequences, including noisy key poses, are iteratively denoised.
@@ 2.1 Human Motion Synthesis
Advancements in data-driven approaches have led to appealing breakthroughs in human motion synthesis. Beyond achieving high-quality and realistic motion outputs, many applications are designed to generate motions adaptable to various conditions, such as action categories [15, 58, 71, 8, 25], text [4, 17, 18, 29, 88, 87, 16, 71, 91, 59, 46, 60, 19], audio [38, 14, 1, 68, 72, 44, 63], interacting objects [83, 12, 43, 95, 6, 37, 70, 3, 42], and scenes [69, 31, 78, 30, 45, 39, 96, 28, 86, 7, 26, 79, 51, 50]. Recently, there has been growing interest in trajectory-based controllable motion generation. For instance, PriorMDM [67] fine-tuned an existing motion diffusion model MDM [71] to support end-effector trajectory control. GMD [67] incorporated classifier-free guidance into a diffusion model to enable root joint-based controllable generation. OmniControl [81] expanded this capability to arbitrary joint trajectories using a combination of ControlNet [90] and inference-time guidance. The recent ControlMM [61] approach, based on masked motion modeling, further enables multi-joint trajectory-based controllable motion generation.
@@ 3.1 Scene Encoding
The global scene features represent the overall spatial context of the entire scene as an occupancy voxel grid at a coarse resolution of $\mathbf{c}_{g}\in\{0,1\}^{d_{x}\times d_{y}\times d_{z}}$ with 0.1m per voxel, where a value of 1 indicates an occupied voxel and 0 represents an unoccupied space. The center and orientation of this grid are initialized by the root position and orientation in the first frame. To effectively encode these global scene features, we employ a Vision Transformer (ViT) model, following the approach in [31]. This global scene encoder processes inputs with dimensions 48×24×4848$\times$ 24$\times$ 4848 × 24 × 48 and results in a 512-dimensional feature vector. The global scene encoder is jointly trained with the rest of the framework.
@@ 3.2 Scene-aware Motion In-betweening
[…]
@@ 3.2.1 Handling Noisy Keyframes
In real-world scenarios, when the keyframe poses are captured by less accurate mocap sensors or extracted from videos, directly using noisy keyposes for motion inbetweening would compromise the output motion quality. Fortunately, diffusion models inherently add noise to motion data during the forward process, and the models learn to remove this noise during the reverse process. As suggested by MotionMix [25] and Rohm [92], the noise in imperfect motions could be analogous to the added noise in the diffusion process at certain timesteps. Motivated by this insight, we divide the diffusion/sampling timesteps into two ranges: $[T,T^{*}+1]$ and $[T^{*},1]$. From early diffusion timestep $T$ to $T^{*}+1$, the prediction of $\mathbf{x}_{0}$ is guided by noisy keyposes $\mathbf{s}^{\text{noisy}}$ for keyframe alignment, while for the remaining timesteps ($T^{*}$ to 111), both keyframes and in-betweening frames are iteratively denoised by $\mathcal{D}_{\theta}$.
@@ 4 Experiments
[…]
@@ 4.1 Evaluation on Noise-free TRUMANS
We compared our model to ablated variants that exclude global scene features $\mathbf{c}_{g}$, local scene features $\mathbf{c}_{l}$, or both (i.e.,“w/o scene awareness“). We also evaluated performance on close-proximity interaction frames where the minimal distance between human and scene is less than 8 cm, with results shown in Tab. 2. From Tab. 1 and Tab. 2, we observed that scene-awareness effectively reduces the collision ratio from 0.237 to 0.162 in interaction frames and enhances alignment with ground-truth motions by over 50% (MJPE from 0.059 to 0.023 m). This observation is consistent with the visualized generation results in Fig. 3. Both global and local scene features enhance motion quality and alignment, with global features playing more prominent roles. Refer to the supplementary video for more results.
@@ 5 Conclusions
Acknowledgements
We sincerely thank the Computational Imaging team at Snap Research NYC, Snap Inc., for their generous support and valuable collaboration. This work was partly supported by IITP grant funded by the Korea government (MSIT) [NO.RS-2021-II211343, Artificial Intelligence Graduate School Program (Seoul National University)] and BK21 FOUR program of the Education and Research Program for Future ICT Pioneers, Seoul National University in 2025.
Supplementary Material
@@ 8.2 Refinement Stage
9 Additional Results
9.1 Evaluation across Multiple Seeds
Since our model is generative, we repeat our major experiments in Table 1 and Table 3 in the main paper across 20 different random seeds. We report their mean value, with 95% statistical confidence interval in Table 9. The observed variance is marginal, demonstrating the stability of our method.
Configurations FID $\downarrow$ Jerk ($m/s^{3}$) $\downarrow$ MJPE ($m$) $\downarrow$ Collision Ratio $\downarrow$
Tab.1 (ours) 0.123 0.194 0.023 0.113
   + 20 runs ${0.123}^{\pm{0.001}}$ ${0.193}^{\pm{0.002}}$ ${0.023}^{\pm{0.001}}$ ${0.114}^{\pm{0.002}}$
Tab.3 (ours) 0.118 0.198 0.012 0.108
   + 20 runs ${0.118}^{\pm{0.001}}$ ${0.198}^{\pm{0.003}}$ ${0.012}^{\pm{0.001}}$ ${0.109}^{\pm{0.002}}$
Table 9: Evaluation across multiple random seeds. We report the mean and 95% confidence intervals for key metrics over 20 runs.
9.2 Integration with Frame-Based HSI
To further explore the applicability of our method, we integrate our module with a semantic keyframe generation approach. Specifically, we generated multiple sparse keyframes (colored in blue) using COINS [95] to provide semantic cues in various scenes, then applied our model to synthesize the complete motion sequence. The Figure 8 show our method generates coherent and plausible motions despite the semantic sparsity of the input keyframes.
Figure 8: Integration with semantically generated keyframes. Our model produces plausible motions from sparse, semantic keyframes.
9.3 Long-Term Keyframe Interval
We also additionally provide an example with a 4-second keyframe interval, where only the start and end frames are given. As shown in Figure 9, the synthesized motion successfully navigates complex scenes with large obstacles, demonstrating our model successfully handles a long motion sequence that navigates around large obstacles.
Figure 9: Result with a long-term keyframe interval. The model synthesizes long-horizon motion while avoiding large obstacles.
9.4 Discussion of Failure Cases
While our method generalizes well to unseen configurations and scene geometries beyond the training data, we acknowledge certain failure cases. First, failure can occur in rare human-scene interaction patterns such as “squeezing through a narrow passage,” where the required motion rarely observed in training. Second, performance degrades in real-world scenes with highly complex or noisy geometric reconstructions, where subtle spatial constraints may not be fully captured by scene encoding. Figure 10 illustrates representative failure cases.
Figure 10: Failure cases. (Left) Unseen interaction pattern (e.g., squeezing through narrow space). (Right) Real-world scene with noisy or complex geometry.